I have an algebra review and forgot how to solve these types of equations..
3n(n-4)(2n+7)=0

Answer by stanbon(75887)

I have an algebra review and forgot how to solve these types of equations..
3n(n-4)(2n+7)=0
------
Since the product is zero, one or more of the factors must be zero.
n = 0 or n-4 = 0 or 2n+7 = 0
-----
n = 0 or n = 4 or n = -7/2
==================================
cheers,
Stan H.
===============